OpenStudy (anonymous):

Please help! A 30.0-V potential difference is applied across a parallel combination of a 60.0-ohm and a 10.0-ohm resistor. What is the current in the 10.0-ohm resistor? Show your work.

OpenStudy (anonymous):

Medals will of course be given once I know it's the right answer :)

OpenStudy (roadjester):

 Is this the image?

OpenStudy (anonymous):

No image came with it

OpenStudy (roadjester):

Ok, so ASSUMING that is the image. The voltage difference across both resistors is the same. So just apply ohm's law.

OpenStudy (anonymous):

I = v/r? I = 30/10? I = 3amps?

OpenStudy (anonymous):

Is that right? If no please help me with the basic equation.

OpenStudy (roadjester):

yes

OpenStudy (anonymous):

Ok thanks!! Once it gets graded, and if it is right, I'll give the medal. :)

OpenStudy (roadjester):

Ok sure

OpenStudy (anonymous):

Wait! to figure out the current you have to find the combined resistance before applying ohms law. with only 2 resistors the formula is easy.\[\frac{R1*R2 }{ R1 +R2 }\] which I calculate to be 8.571 ohms. Take 30v / 8.571 = 3.5 amps